I found a very similar problem to this one but when I went through all the steps the answer came out negative and Im pretty sure you cant have a negative speed?
Alvin paddled for 4 hours with a 6km/h current to reach a campsite. The return trip against the current took 10 hours. Find the speed of Alvin's canoe in still water.
: I found a very similar problem to this one but when I went through all the steps the answer came out negative and Im pretty sure you cant have a negative speed?
Alvin paddled for 4 hours with a 6km/h current to reach a campsite. The return trip against the current took 10 hours. Find the speed of Alvin's canoe in still water.

Answer by edjones(2120)

st=d, s=speed in still water, t=time, d=distance
4(s+6)=d
10(s-6)=d
4(s+6)=10(s-6)
4s+24=10s-60
84=6s
s=14 km/h in still water.
.
Ed
Question 149180: I found a very similar problem to this one but when I went through all the steps the answer came out negative and Im pretty sure you cant have a negative speed?
Alvin paddled for 4 hours with a 6km/h current to reach a campsite. The return trip against the current took 10 hours. Find the speed of Alvin's canoe in still water.
: I found a very similar problem to this one but when I went through all the steps the answer came out negative and Im pretty sure you cant have a negative speed?
Alvin paddled for 4 hours with a 6km/h current to reach a campsite. The return trip against the current took 10 hours. Find the speed of Alvin's canoe in still water.

Answer by josmiceli(1827)

Let s = the speed of Alvin's canoe
Let d = the length of the trip
Going with the current:
d = (s + 6)*t
t = 4 hrs is given
d = (s + 6)*4 km
Going against the current:
d = (s - 6)*10 km
Make the 2 expressions for d equal to each other
(s + 6)*4 = (s - 6)*10
4s + 24 = 10s - 60
6s = 84
s = 14
The speed of the boat in still water is 14 km/hr
check:
d = (s + 6)*4
d = (14 + 6)*4
d = 80 km
and
d = (s - 6)*10
d = (14 - 6)*10
d = 80 km
OK